An experiment started with 600 bacteria. They double in number every hour. Find the number of bacteria after 4 hours.
Answer by nerdybill(7384)

An experiment started with 600 bacteria. They double in number every hour. Find the number of bacteria after 4 hours.
1200 = 600e^(1r)
1200/600 = e^(r)
2 = e^(r)
ln(2) = r
.
general equation:
A = 600e^(tln(2))
for 4 hours:
A = 600e^(4ln(2))
A = 600e^(2.77259)
A = 9600 bateria